OpenStudy (anonymous):

Medal A cafeteria sold 156 cartons of milk in 1 day. Each carton contained 1 c of milk. How many quarts of milk were sold that day? ? qt

OpenStudy (anonymous):

your hot

OpenStudy (anonymous):

who is this?

OpenStudy (anonymous):

he says that to everyone

OpenStudy (elonasushchik):

Huh..

OpenStudy (anonymous):

oh lol

OpenStudy (anonymous):

do u mean 1 g of milk?

OpenStudy (elonasushchik):

He dosnt know what he's talking about ...

OpenStudy (elonasushchik):

Not u skate, but smarty..

OpenStudy (anonymous):

no your the first girl shut up @skate-4-life270

OpenStudy (anonymous):

no 1 day

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i mean the other 1

OpenStudy (anonymous):

no its what it says in the question

OpenStudy (anonymous):

maybe it was a typo

geerky42 (geerky42):

My hot what? @Smartypants123

OpenStudy (anonymous):

i dont know maybe it is

OpenStudy (elonasushchik):

1 US Quart=4 US Cups so...

OpenStudy (anonymous):

i really dont understand this question

OpenStudy (jhannybean):

You have 156 cartons of milk a day 1 carton = 1 cup C1 qt of milk = 4 cups

OpenStudy (jhannybean):

So \[156~\text{cartons} \cdot \frac{1 ~\text{cup}}{1~\text{carton}} \cdot \frac{1~\text{quart}}{4~\text{cups}}\]

OpenStudy (jhannybean):

You are being asked to convert cartons of milk \(\to\) quarts of milk. That is all you have to do.

OpenStudy (anonymous):

ok so u cross multiply

OpenStudy (jhannybean):

Do you know how to solve by dimensional analysis?

OpenStudy (anonymous):

no

OpenStudy (jhannybean):

What I wanted to do was get from the amount of cartons to the number off cups given (in the problem) then use a known conversion factor to get from the number of cups in a quart of milk. You will see that the units cancel out, \[156~\cancel{\text{cartons}} \cdot \frac{1 ~\cancel{\text{cup}}}{1~\cancel{\text{carton}}} \cdot \frac{1~\text{quart}}{4~\cancel{\text{cups}}}\] And we will be left with the number of quarts.

OpenStudy (anonymous):

39

OpenStudy (jhannybean):

Now we just multiply across, \[\frac{156(1)(1~\text{quart})}{(1)(4)(1)}\]

OpenStudy (anonymous):

624

OpenStudy (jhannybean):

and yes, you end up with 39 quarts.

OpenStudy (jhannybean):

i do not know how you are getting these numbers. Reread my posts for a better understanding.

OpenStudy (anonymous):

156 * 1*1/1*4 = 624

OpenStudy (jhannybean):

That is not correct.. 156/4 \(\ne\) 624.

OpenStudy (anonymous):

so the answer is 39 qt?

OpenStudy (jhannybean):

Yes, you are right.

OpenStudy (anonymous):

thank you :)

OpenStudy (jhannybean):

No problem. I think where you were getting 624 from was 156 * 4. but we were dividing instead of multiplying.

OpenStudy (anonymous):

ya sorry

OpenStudy (jhannybean):

No problem:)